# A Survey on Open Information Extraction from Rule-based Model to Large Language Model

### v6

@@ Abstract
Open Information Extraction (OpenIE) represents a crucial NLP task aimed at deriving structured information from unstructured text, unrestricted by relation type or domain. This survey paper provides an overview of OpenIE technologies spanning from 2007 to 2024, emphasizing a chronological perspective absent in prior surveys. It examines the evolution of task settings in OpenIE to align with the advances in recent technologies. The paper categorizes OpenIE approaches into rule-based, neural, and pre-trained large language models, discussing each within a chronological framework. Additionally, it highlights prevalent datasets and evaluation metrics currently in use. Building on this extensive review, the paper outlines potential future directions in terms of datasets, information sources, output formats, methodologies, and evaluation metrics.
@@ A Survey on Open Information Extraction from Rule-based Model to Large Language Model
Pai Liu1,2,thanks: These authors contribute to this work equally. Names are ordered randomly., Wenyang Gao1,11footnotemark: 1, Wenjie Dong3,11footnotemark: 1, Lin Ai4,11footnotemark: 1, Ziwei Gong4,11footnotemark: 1, Songfang Huang5, Zongsheng Li6, Ehsan Hoque 2 Julia Hirschberg4, Yue Zhang1,thanks: The corresponding author, 1Westlake University, 2University of Rochester, 3Zhejiang University 4Columbia University, 5Alibaba DAMO Academy, 6Northeastern University zhangyue@westlake.edu.cn
@@ 1 Introduction
Open Information Extraction (OpenIE) aims to extract structured information from unstructured text sources Niklaus et al. (2018), typically outputting relationships as triplets $(arg_{1},rel,arg_{2})$. Unlike traditional IE, which relies on predefined categories to identify relationships, OpenIE operates without such constraints, as illustrated in Figure 1, enabling the extraction of diverse and unforeseen relations. This flexibility makes OpenIE especially valuable for rapidly evolving Natural Language Processing (NLP) tasks such as question answering, search engines, and knowledge graph completion Han et al. (2020), as well as for handling large-scale and dynamic data sources like web data. Its capability to adapt to new contexts and extract a broader array of information without extensive manual input not only enhances scalability but also broadens the applicative horizon, placing OpenIE at the forefront of NLP research and attracting significant interest for its potential.
Figure 1: Comparison of OpenIE and traditional relation extraction.
Figure 2: Chronological overview of Open IE methods.
@@
Since its inception in 2007, the field of OpenIE has embodied the spirit of relentless innovation within NLP. Initially utilizing basic linguistic tools, OpenIE models have progressively integrated more complex syntactic and semantic features, while preserving the intuitive task of directly extracting relational triplets from text. The advent of neural models in 2019 marks a game-changer for OpenIE research, with systems employing Transformer-based architectures like BERT Devlin et al. (2019) to significantly enhance feature extraction capabilities. To accommodate the technological shift, a variety of methodologies and task settings have evolved within diversified OpenIE approaches.
@@
The emergence of Large Language Models (LLMs) in 2023 has marked another revolutionary phase, steering OpenIE toward a generative method of information extraction. The robust generalization abilities of these models not only advance the technical prowess of OpenIE systems but also facilitate a convergence of methodologies and task settings – revisiting the original, straightforward text $\rightarrow$ relational triplet format. This transition also fosters potential integration with traditional IE tasks, pointing toward a promising future where extraction tasks are tackled through a unified, multi-task approach. With these advancements, OpenIE is poised not only to adapt to evolving technologies but also to influence the trajectory of other NLP downstream tasks.
The field of OpenIE has been marked by continuous innovation and evolution, yet there is a notable gap in the literature. Previous surveys largely focus on pre-LLM era models or limit their scope to methodological insights Gamallo (2014); Vo and Bagheri (2018); Zouaq et al. (2017); Glauber and Claro (2018); Niklaus et al. (2018); Zhou et al. (2022). Moreover, while recent studies Xu et al. (2023b) delve into information extraction in the LLM era, they largely bypass OpenIE, concentrating instead on traditional IE tasks. This oversight highlights the urgent need for a comprehensive, chronological review that captures the full trajectory of OpenIE’s development alongside advancements in broader NLP technologies. Undertaking such an in-depth exploration not only provide a macroscopic view of the field’s historical progression but also illuminate future possibilities and opportunities within OpenIE, giving scholars and researchers a clearer roadmap for advancing this crucial area of study.
The remainder of this review paper is organized as follows. Section 2 details the primary OpenIE task setting and its variants. Section 3 and 4 cover popular datasets and evaluation metrics, respectively, followed by a chronological review of OpenIE models in Section 5, illustrated by Fig.2. The sources of information are thoroughly presented in Section 6.1, while Section 6.2 explores current limitations and future prospects of OpenIE.
@@ 2 Task Settings
Figure 3: An overview of workflow processes in OpenIE task settings.
We categorize OpenIE task settings into three groups: Open Relation Triplet Extraction (ORTE), Open Relation Span Extraction (ORSE) and Open relation clustering (ORC). ORTE is the classic task setting, while ORSE and ORC settings are variations developed to cater to diverse models with the advancement of NLP techniques. For all three task settings, the openness is shown in the absence of restraints on relation types. Figure 2 illustrates an overall review of the chronological development of task settings. Figure 3 depicts the workflow for each task setting.
@@
Banko et al. (2007) initially defines open information extraction as an unsupervised task that automatically extracts $(entity_{1},relation,entity_{2})$ triplets from a vast corpus of unstructured web text, where $entity_{1}$, $entity_{2}$ and $relation$ consist of selected words from input sentences. This task setting, irrespective of the learning method or the forms of input and output, represents the most idealized configuration.
@@
Open relation clustering (ORC), also widely known as open relation extraction, clusters relation instances $(h,t,s)$, where $h$ and $t$ denote head entity and tail entity respectively, and $s$ denotes the sentence corresponding to two entities. Different from the ORTE, ORC does not extract entities from text but uses the whole text to represent the relation between two entities. Clustering similar relations is a step forward in labeling specific relations to each relation instance. These task settings outlined above are distinctly characterized by era-specific traits and methodologies, further discussed in Section 5.
@@ 3 Datasets
Previous surveys Niklaus et al. (2018); Ali et al. (2019) conclude with models of the first two generations and the datasets they used. We exclude those small-scale datasets and some seldom-used datasets. Table 1 lists some popular and promising OpenIE datasets grouped by their creating methods.
Question Answering (QA) derived datasets are converted from other crowd-sourced QA datasets. OIE2016 Stanovsky and Dagan (2016) is one of the most popular OpenIE benchmarks, which leverages QA-SRL He et al. (2015) annotations. Additional datasets extend from OIE2016, such as AW-OIE Stanovsky et al. (2018), Re-OIE2016 Zhan and Zhao (2020) and CaRB Bhardwaj et al. (2019). LSOIE Solawetz and Larson (2021), is created by converting the QA-SRL 2.0 dataset FitzGerald et al. (2018) to a large-scale OpenIE dataset, which claims to be 20 times larger than the next largest human-annotated OpenIE dataset.
@@
Instruction-based datasets transform IE tasks into tasks requiring instruction-following, thus harnessing the capabilities of LLMs. Strategies include integrating existing IE datasets into a unified-format Wang et al. (2023a); Lu et al. (2022), and deriving others from Wikidata and Wikipedia such as InstructOpenWiki Lu et al. (2023), InstructIE Gui et al. (2023), and Wikidata-OIE Wang et al. (2022b).
Overall, KB derived datasets are mostly used in open relation clustering task settings, illustrated in Section 5.4, whereas QA derived and crowd-sourced datasets are usually used in open relational triplet extraction (Section 5.2) and open relation span extraction task settings (Section 5.3). We provide more detailed descriptions in Appendix D.
Dataset #Tuple Domain
@@
OIE2016 (2016) 10,359 Wiki, Newswire
Re-OIE2016 (2020) NR Wiki, Newswire
CaRB (2019) NR Wiki, Newswire
AW-OIE (2018) 17,165 Wiki, Wikinews
LSOIE-wiki (2021) 56,662 Wiki, Wikinews
LSOIE-sci (2021) 97,550 Science
@@
WiRe57 (2019) 343 Wiki, Newswire
$\textrm{SAOKE}^{{zh}}$ (2018) NR Baidu Baike
$\textrm{BenchIE}^{{en}}$ (2021) 136,357 Wiki, Newswire
$\textrm{BenchIE}^{{de}}$ (2021) 82,260 Wiki, Newswire
$\textrm{BenchIE}^{{zh}}$ (2021) 5,318 Wiki, Newswire
@@
NYT-FB (2008; 2008; 2009; 2011) 39,000 NYT, Freebase
TACRED (2017) 119,474 TAC KBP
FewRel (2018) 70,000 Wiki, Wikidata
T-REx (2018) 11M Wiki, Wikidata
$\textrm{COER}^{{zh}}$ (2018) 1M Baidu Baike, Chinese news
@@
InstructOpenWiki (2023) 19M Wiki, Wikidata
Wikidata-OIE (2022b) 27M Wiki, Wikidata
Table 1: Statistics of popular OpenIE datasets. "NR" stands for "Not Reported". Non-English datasets are indicated with superscripts.
@@ 4 Evaluation
Task Setting Evaluation Metrics
ORTE
Recall, AUC, F1
ORSE F1
ORC ARI, $B^{3}$, V-measure
Table 2: Core evaluation metrics of each task setting.
Evaluation metrics for OpenIE models vary by task setting. In the open relational triplet and relation span extraction settings (Sections 5.2 and 5.3), models are assessed using precision, recall, F1 score, and AUC, potentially employing various scoring functions. In the open relation clustering setting (Section 5.4), performance is evaluated using $B^{3}$ Bagga and Baldwin (1998), V-measure Rosenberg and Hirschberg (2007), and ARI Hubert and Arabie (1985).
In addition to standard metrics, various methods employ additional metrics, typically categorized into token-level and fact-level scorers. Token-level scorers focus on individual tokens to ensure precision and semantic accuracy, accommodating linguistic variability Stanovsky and Dagan (2016), enhancing conciseness Léchelle et al. (2019), and adapting to complex model outputs like those from LLMs Han et al. (2023). Fact-level scorers assess the informational faithfulness of extractions to ensure reliable knowledge extraction, validating semantic and information integrity Sun et al. (2018); Gashteovski et al. (2021); Li et al. (2023a) to enhance OpenIE evaluations comprehensively. Further details are discussed in Appendix E.
5 Methodologies
Research on OpenIE dates back to 2007, with the first generation of OpenIE models exemplified by TextRunner Banko et al. (2007), WOE Wu and Weld (2010) and ReVerb Fader et al. (2011), which use shallow linguistic features such as part-of-speech (POS) tags and noun phrase (NP) chunk features. The second generation of OpenIE models, represented by OLLIE Schmitz et al. (2012), ClausIE Del Corro and Gemulla (2013), SRL-IE Christensen et al. (2010) and OpenIE4 Mausam (2016), incorporates deep linguistic features alongside shallow syntactic features. The third generation of OpenIE models, to be discussed in detail, benefits significantly from the advent of neural models such as Transformers Vaswani et al. (2017), notably BERT Devlin et al. (2019), and extensively employ pre-training models for feature extraction. The most recent, fourth-generation has emerged with the advancements in Large Language Models (LLMs), such as GPT models OpenAI (2024, 2023) and Llama 2 Touvron et al. (2023), leading to generative IE methods. These innovations highlight a shift toward universal formats and multi-task models, indicating new trends in the future of OpenIE. In Fig.2, we present OpenIE methodologies in a chronological view and categorize them according to task settings instead of the approaches. More details about model implementation is provided in Appendix.B
@@ 5 Methodologies
OpenIE System F1 AUC F1 AUC F1 AUC ARI $B^{3}$ V ARI $B^{3}$ V
Pre-Neural (ORTE) Ollie Schmitz et al. (2012) 38.6 20.2 49.5 31.3 41.1 22.4 - - - - - -
@@
Neural Era (ORTE) OpenIE6 Kolluru et al. (2020a) - - - - 52.7 33.7 - - - - - -
IMoJIE Kolluru et al. (2020b) - - - - 53.5 33.3 - - - - - -
Neural Era (ORSE) $\textrm{Multi}^{2}\textrm{OIE}$ Ro et al. (2020) - - 83.9 74.6 52.3 32.6 - - - - - -
OIE@OIA Wang et al. (2022e) 71.6 54.3 85.3 76.9 51.1 33.9 - - - - - -
Neural Era (ORC) SelfORE Hu et al. (2020) - - - - - - 64.7 67.8 78.3 44.7 54.1 61.9
MatchPrompt Wang et al. (2022c) - - - - - - 66.5 72.3 82.2 75.3 83.0 84.5
LLM Era IELM GPT-2XL Wang et al. (2022b) - - 35.0 - 22.7 - - - - - - -
GPT-3.5-Turbo ICL Ling et al. (2023) 65.1 - 67.9 - 52.1 - - - - - - -
Table 3: Performance of selected representative OpenIE models. Only report F1 scores for $B^{3}$ and V. Complete performance details in the Appendix A.

### v7

@@ Abstract
Open Information Extraction (OpenIE) represents a crucial NLP task aimed at deriving structured information from unstructured text, unrestricted by relation type or domain. This survey paper provides an overview of OpenIE technologies spanning from 2007 to 2024, emphasizing a chronological perspective absent in prior surveys. It examines the evolution of task settings in OpenIE to align with the advances in recent technologies. The paper categorizes OpenIE approaches into rule-based, neural, and pre-trained large language models, discussing each within a chronological framework. Additionally, it highlights prevalent datasets and evaluation metrics currently in use. Building on this extensive review, this paper systematically reviews the evolution of task settings, data, evaluation metrics, and methodologies in the era of large language models, highlighting their mutual influence, comparing their capabilities, and examining their implications for open challenges and future research directions.
@@ A Survey on Open Information Extraction from Rule-based Model to Large Language Model
Pai Liu1,2,thanks: These authors contributed equally to this work., Wenyang Gao1,11footnotemark: 1, Wenjie Dong3,11footnotemark: 1, Lin Ai4,11footnotemark: 1, Ziwei Gong4,11footnotemark: 1, Songfang Huang5, Zongsheng Li6, Ehsan Hoque 2 Julia Hirschberg4, Yue Zhang1,thanks: The corresponding author., 1Westlake University, 2University of Rochester, 3Zhejiang University 4Columbia University, 5Alibaba DAMO Academy, 6Northeastern University zhangyue@westlake.edu.cn
@@ 1 Introduction
Open Information Extraction (OpenIE) aims to extract structured information from unstructured text sources Niklaus et al. (2018), typically outputting relationships as triplets $(arg_{1},rel,arg_{2})$. As illustrated in Figure 1, unlike standard IE, which relies on predefined categories to identify relationships, OpenIE operates without such constraints, enabling the extraction of diverse and unforeseen relations. This flexibility makes OpenIE especially valuable for rapidly evolving Natural Language Processing (NLP) tasks such as question answering, search engines, and knowledge graph completion Han et al. (2020), as well as for handling large-scale and dynamic data sources like web data.
Figure 1: Comparison of OpenIE and standard relation extraction.
@@
Since its inception in 2007, the field of OpenIE has witnessed continual advancements. Initially utilizing basic linguistic tools, OpenIE models have progressively integrated more complex syntactic and semantic features, while preserving the intuitive task of directly extracting relational triplets from text. The advent of neural models in 2019 marks a paradigm shift for OpenIE research, where systems employing Transformer-based architectures like BERT Devlin et al. (2019) significantly enhance feature extraction capabilities. To accommodate the technological shift, a variety of methods and task settings have evolved within diversified OpenIE approaches.
@@
The emergence of Large Language Models (LLMs) in 2023 has marked another revolutionary phase, steering OpenIE toward generative information extraction. The robust generalization abilities of these models not only advance the technical prowess of OpenIE systems but also facilitate the convergence of methods and task settings – revisiting the original, straightforward text $\rightarrow$ relational triplet format. This transition fosters potential integration with standard IE tasks, pointing toward a promising future where extraction tasks are tackled through a unified, multi-task approach.
As a result, there has been a decline in OpenIE research in the LLM era. Is OpenIE research going to its end? How can traditional OpenIE research inspire IE research in the LLM era? Previous surveys largely focus on pre-LLM era models or limit their scope to methodological insights Gamallo (2014); Vo and Bagheri (2018); Zouaq et al. (2017); Glauber and Claro (2018); Niklaus et al. (2018); Zhou et al. (2022). While recent studies Xu et al. (2023b) delve into information extraction in the LLM era, they largely bypass OpenIE, concentrating instead on standard IE tasks. We aim to bridge this gap by providing a holistic survey of the OpenIE field from a chronological view, addressing the two research questions above.
From a chronological perspective, we summarize all task settings (Section 2), data (Section 3), evaluation metrics (Section 4), and mainstream methods (Section 5) before and after LLM era. We use a single table to summarize mainstream methods and results from different periods. We emphasize the co-evolution between models and task settings, and the various sources of information used to address Open challenges. Based on this, we compare the ideas and relative strengths and weaknesses of large models and traditional models (Section 6.1), review the impact of large language models on open information extraction and traditional methods (Section 6.2), and explore future directions (Section 6.3).
@@ 2 Task Settings
Figure 2: An overview of workflow processes in OpenIE task settings. ORTE aims to extract all n-ary relation tuples in the input, . ORSE finds relational spans according to previously extracted subjects and Objects. ORC pairs the input sentence with different subjects and objects within the sentence to form relation instances, relation instances are iteratively optimized in a supervised, unsupervised, or semi-supervised manner, and after the representations converge, clustering is performed. Objects of the same color indicate that they belong to the same relation cluster. To facilitate observation, we have bolded the borders of the three examples in the figure. Each circle represents the relation instances clustered into the same class by the clustering algorithm.
We categorize OpenIE task settings into three groups: Open Relation Triplet Extraction (ORTE), Open Relation Span Extraction (ORSE) and Open relation clustering (ORC). ORTE is the classic task setting, while ORSE and ORC settings are variations developed to cater to diverse models with the advancement of NLP techniques. For all three task settings, openness is shown in the absence of restraints on relation types. Figure 2 depicts the workflow for each task setting.
@@
Banko et al. (2007) initially defines open information extraction as an unsupervised task that automatically extracts $(entity_{1},relation,entity_{2})$ triplets from a vast corpus of unstructured web text, where $entity_{1}$, $entity_{2}$ and $relation$ consist of selected words from input sentences. Although the term triplet is more commonly used, the actual extraction tasks are not always limited to triplets and can involve more diverse n-ary relations, such as condition, temporal information, etc. This task setting, irrespective of the learning method or the forms of input and output, represents the most idealized configuration.
@@
Open relation clustering (ORC), also known as open relation extraction, clusters relation instances $(h,t,s)$, where $h$ and $t$ denote head entity and tail entity respectively, and $s$ denotes the sentence corresponding to two entities. Different from the ORTE, ORC does not extract relation from text but uses text between two entities to represent the relation. Clustering similar relations is a step forward in labeling specific relations to each relation instance. These task settings outlined above are distinctly characterized by era-specific traits and methodologies, further discussed in Section 5.
Summarized in Table 2, ORTE efficiently identifies all relations in one step, reducing error and ensuring consistency, but struggles with complex implementation, overlapping relations, and large data requirements. ORSE simplifies individual models, handling nested relations by focusing on specific subtasks, but is prone to error propagation and is computationally costly. ORC clusters diverse relation expressions to aid generalization, but faces scalability challenges due to intensive similarity computations and potential loss of specificity.
@@ 3 Datasets
Table 1 lists some popular and promising OpenIE datasets grouped by their creating methods.
Question Answering (QA) derived datasets are converted from other crowd-sourced QA datasets. OIE2016 Stanovsky and Dagan (2016), one of the most popular OpenIE benchmarks, leverages QA-SRL He et al. (2015) annotations. Additional datasets extend from OIE2016, such as AW-OIE Stanovsky et al. (2018), Re-OIE2016 Zhan and Zhao (2020) and CaRB Bhardwaj et al. (2019). LSOIE Solawetz and Larson (2021), is created by converting the QA-SRL 2.0 dataset FitzGerald et al. (2018) to a large-scale OpenIE dataset, which claims to be 20 times larger than the next largest human-annotated OpenIE dataset.
@@
Instruction-based datasets transform IE tasks into instruction-following tasks to harness the capabilities of LLMs. Strategies include integrating existing IE datasets into a unified-format Wang et al. (2023a); Lu et al. (2022), and deriving others from Wikidata and Wikipedia such as InstructOpenWiki Lu et al. (2023), InstructIE Gui et al. (2023), and Wikidata-OIE Wang et al. (2022b).
Overall, KB derived datasets are mostly used in ORC task settings, whereas QA derived, crowd-sourced, and instruction-based datasets are usually used in ORTE and ORSE task settings. We provide more detailed descriptions in Appendix C.
Dataset #Tuple Domain Task
@@
OIE2016 (2016) 10,359 Wiki, Newswire ORTE/ORSE
Re-OIE2016 (2020) NR Wiki, Newswire ORTE/ORSE
CaRB (2019) NR Wiki, Newswire ORTE/ORSE
AW-OIE (2018) 17,165 Wiki, Wikinews ORTE/ORSE
LSOIE-wiki (2021) 56,662 Wiki, Wikinews ORTE/ORSE
LSOIE-sci (2021) 97,550 Science ORTE/ORSE
@@
WiRe57 (2019) 343 Wiki, Newswire ORTE/ORSE
$\textrm{SAOKE}^{{zh}}$ (2018) NR Baidu Baike ORC
$\textrm{BenchIE}^{{en}}$ (2021) 136,357 Wiki, Newswire ORTE/ORSE
$\textrm{BenchIE}^{{de}}$ (2021) 82,260 Wiki, Newswire ORTE/ORSE
$\textrm{BenchIE}^{{zh}}$ (2021) 5,318 Wiki, Newswire ORTE/ORSE
@@
NYT-FB (2008; 2008; 2009; 2011) 39,000 NYT, Freebase ORC
TACRED (2017) 119,474 TAC KBP ORC
FewRel (2018) 70,000 Wiki, Wikidata ORC
T-REx (2018) 11M Wiki, Wikidata ORC
$\textrm{COER}^{{zh}}$ (2018) 1M Baidu Baike, Chinese news ORC
@@
InstructOpenWiki (2023) 19M Wiki, Wikidata ORTE/ORSE
Wikidata-OIE (2022b) 27M Wiki, Wikidata ORTE/ORSE
Table 1: Statistics of popular OpenIE datasets. "NR" stands for "Not Reported". Non-English datasets are indicated with superscripts. The Task column indicates the types of tasks the data can be used for.
@@ 4 Evaluation
Task Setting Datasets Evaluation Metrics Advantages / Disadvantages
ORTE OIE2016 Re-OIE2016 CaRB Precision Recall F1 AUC Advantages: Efficiency; Reduced Error Propagation Consistency Disadvantages: Difficulty in Handling Overlapping Relations; Handling complex relations
ORSE OIE2016 Re-OIE2016 CaRB Precision Recall F1 AUC Advantages: Control Over Intermediate Steps; Handling of Overlapping Relations; Modular approach allows for specialized improvements; More Explainable Disadvantages: Error propagation; Increased Computational Cost; Complexity in Integration
ORC NYT-FB FewRel TACRED T-REx B3 V-measure ARI Advantages: Flexibility; Generalization; Lower dependency on annotated data; Knowledge Organization Disadvantages: Lack of explicit relation labels, Ambiguity in clusters; Scalability Issues; Loss of Specificity
Table 2: Overview of task settings in Open Information Extraction (OpenIE), detailing frequently used datasets, typical evaluation metrics, and a summary of key advantages and disadvantages associated with each task.
Evaluation metrics for OpenIE models vary by task setting. In the ORTE and ORSE settings, models are assessed using precision, recall, F1 score, and AUC, potentially employing various scoring functions. In the ORC setting, performance is evaluated using $B^{3}$ Bagga and Baldwin (1998), V-measure Rosenberg and Hirschberg (2007), and ARI Hubert and Arabie (1985).
To compare the extracted and golden triplets, various datasets employ different matching strategies, typically categorized into token-level and fact-level scorers. Token-level scorers focus on individual tokens to ensure precision and semantic accuracy, accommodating linguistic variability Stanovsky and Dagan (2016), enhancing conciseness Léchelle et al. (2019), and adapting to complex model outputs like those from LLMs Han et al. (2023). Fact-level scorers assess the informational faithfulness of extractions to ensure reliable knowledge extraction, validating semantic and information integrity Sun et al. (2018); Gashteovski et al. (2021); Li et al. (2023a) to enhance OpenIE evaluations comprehensively. Further details are discussed in Appendix D.
From the perspective of task formulation, token-level scorers are better suited for open relation span extraction (ORSE), where outputs are succinct, and labeling models in open relation triplet extraction (ORTE), whose outputs are precise tokens derived from the inputs. Conversely, fact-level scorers are more appropriate for generative models in ORTE, particularly LLMs, whose outputs exhibit diversity and necessitate semantic evaluation.
@@
Category Representative Approach F1 AUC F1 AUC F1 AUC ARI $B^{3}$ V ARI $B^{3}$ V
- Ollie Schmitz et al. (2012) 38.6 20.2 49.5 31.3 41.1 22.4 - - - - - -
@@
- OpenIE4 Mausam (2016) 58.8 40.8 68.3 50.9 51.6 29.5 - - - - - -
Pre-Neural (ORTE) 2007 - 2018 - PropS Stanovsky and Dagan (2016) 54.4 32.0 64.2 43.3 31.9 12.6 - - - - - -
SL RnnOIE Stanovsky et al. (2018) 62.0 48.0 - - 49.0 26.1 - - - - - -
SL OpenIE6 Kolluru et al. (2020a) - - - - 52.7 33.7 - - - - - -
Span SpanOIE Zhan and Zhao (2020) 69.4 49.1 77.0 65.8 48.5 - - - - - - -
S2S IMoJIE Kolluru et al. (2020b) - - - - 53.5 33.3 - - - - - -
SL MacroIE Bowen et al. (2021) - - - - 54.8 36.3 - - - - - -
SL DetIELSOIE Vasilkovsky et al. (2022) - - - - 43.0 27.2 - - - - - -
SL DetIEIMoJIE Vasilkovsky et al. (2022) - - - - 52.1 36.7 - - - - - -
Neural Era (ORTE) 2018 - 2022 SL SMiLe-OIE Dong et al. (2022) - - - - 53.8 34.9 - - - - - -
SL $\textrm{Multi}^{2}\textrm{OIE}$ Ro et al. (2020) - - 83.9 74.6 52.3 32.6 - - - - - -
S2S Gen2OIE Kolluru et al. (2022) - - - - 54.4 32.3 - - - - - -
S2S Gen2OIE (label-rescore) - - - - 54.5 38.9 - - - - - -
Graph OIE@OIA Wang et al. (2022d) 71.6 54.3 85.3 76.9 51.1 33.9 - - - - - -
Graph DragonIE Yu et al. (2022) - - - - 55.1 36.4 - - - - - -
Graph ChunkOIE(SaC-OIA-SP) Dong et al. (2023) - - - - 53.6 35.5 - - - - - -
Neural Era (ORSE) 2018 - 2022 Graph ChunkOIE(SaC-CoNLL) - - - - 53.2 34.7 - - - - - -
Semi RSN Wu et al. (2019) - - - - - - 45.3 58.9 70.8 45.9 63.1 64.3
Semi RSN-CV Wu et al. (2019) - - - - - - 54.2 63.8 72.4 - - -
Un SelfORE Hu et al. (2020) - - - - - - 64.7 67.8 78.3 44.7 54.1 61.9
Semi RSN-BERT Zhao et al. (2021) - - - - - - 53.2 70.9 78.1 75.6 83.4 85.9
Semi RoCORE Zhao et al. (2021) - - - - - - 70.9 79.6 86 81.2 86 88.8
Semi OHRE Zhang et al. (2021a) - - - - - - 64.2 70.5 76.7 - - -
Semi MatchPrompt Wang et al. (2022c) - - - - - - 66.5 72.3 82.2 75.3 83.0 84.5
Un PromptORE Genest et al. (2022) - - - - - - 43.4 48.8 71.8 - - -
Semi CaPL Duan et al. (2022) - - - - - - 79.4 81.9 88.9 82.9 87.3 89.8
Neural Era (ORC) 2018 - 2022 Semi ASCORE Zhao et al. (2023) - - - - - - 67.6 73.5 83.5 78.1 78 83.1
0-shot IELM GPT-2XL Wang et al. (2022b) - - 35.0 - 22.7 - - - - - - -
few-shot GPT-3.5-Turbo ICL Ling et al. (2023) 65.1 - 67.9 - 52.1 - - - - - - -
Transition Era to LLM (ORTE) 2022 - few-shot ChatGPT $n$-shot Qi et al. (2023a) - - - - 55.3 - - - - - - -
Table 3: Performance of OpenIE models. For $B^{3}$ and V measures, F1 scores are reported. Rows filled with colors represent models of different task settings:   = ORSE,   = ORSE,   = ORC. SL stands for sequence labeling, S2S denotes sequence to sequence, Span represents a span-based method, Graph signifies a graph-based method, Semi refers to semi-supervised learning, and Un refers to unsupervised learning.
5 A Chronological Review of Mainstream Methods
The research approaches for Open IE have undergone three significant changes along with technological advancements. We categorize these periods into three eras: the pre-neural era, dominated by rule-based and statistic-based methods; the neural model era, primarily based on neural networks; and the LLMs era, characterized by the use of LLMs. Chronologically, we will discuss the key models and methods from each period and explore their connections, as depicted in Figure 3. More details on model implementation provided in Appendix.A
Figure 3: Chronological overview of Open IE methods.